OpenStudy (anonymous):

What is the value of x? 8 cm 14 cm 18 cm 28 cm

OpenStudy (anonymous):

figure

OpenStudy (anonymous):

Similar triangles. The ratio of the sides are equal to each other, only this time, they were jerks for making the length x + 2.... Still, not a problem, (x + 2)/14 = (x + 10)/21 x + 2 = 2/3 * x + 2/3 * 10 1/3 x = 2/3 * 10 - 2 x = 20 -6 x = 14 Double check my arithmetic!

OpenStudy (anonymous):

looks good :D thank you